At the beginning of Year 2, the Redd Company had the following balances in its accounts:

Cash $ 8,000
Inventory 2,000
Common stock 7,500
Retained earnings 2,500


During Year 2, the company experienced the following events:

  1. Purchased inventory that cost $5,500 on account from Ross Company under terms 2/10, n/30. The merchandise was delivered FOB shipping point. Freight costs of $500 were paid in cash.
  2. Returned $350 of the inventory it had purchased because the inventory was damaged in transit. The seller agreed to pay the return freight cost.
  3. Paid the amount due on its account payable to Ross Company within the cash discount period.
  4. Sold inventory that had cost $6,000 for $9,000 on account, under terms 2/10, n/45.
  5. Received merchandise returned from a customer. The merchandise originally cost $500 and was sold to the customer for $800 cash. The customer was paid $800 cash for the returned merchandise.
  6. Delivered goods FOB destination in Event 4. Freight costs of $600 were paid in cash.
  7. Collected the amount due on the account receivable within the discount period.
  8. Took a physical count indicating that $1,700 of inventory was on hand at the end of the accounting period.

b. Record each event in a statements model like the following one. In the Cash Flow column, use OA to designate operating activity, IA for investment activity, FA for financing activity, NC for net change in cash and NA to indicate the element is not affected by the event. The first event is recorded as an example.

REDD COMPANY
Horizontal Statements Model - Year 2
Event Balance Sheet Income Statement Statement of Cash Flows
Assets = Liabilities + Stockholders’ Equity Revenue Expenses = Net Income
Cash + Accounts Receivable + Merchandise Inventory = Accounts Payable + Common Stock + Retained Earnings
Bal. 8,000 + + 2,000 = + 7,500 + 2,500 = NA
1a. + + = + + = 0 NA
1b. (500) + + 500 = + + = (500) OA
2. + + (350) = (350) + + = 0 NA
3. (4,120) + + (1,030) = (1,030) + + =
4a. + 9,000 + = + + 9,000 9,000 = 9,000 0 NA
4b. + + (6,000) = + + (6,000) 6,000 = (6,000) 0 NA
5a. (800) + + = + + (800) (800) = (800) (800) OA
5b. + + 500 = + + 500 (500) = 500 0 OA
6. (600) + + = + + (600) 600 = (600) (600) OA
7. + (9,000) + = + + ( = OA
8. + + 1,030 = + + 1,030 1,030 = 1,030 0 NA
Bal. $10,802 + $0 + $2,150 = $4,120 + $7,500 + $5,452 $9,052 $6,100 = $2,952 $6,922 NC

can you please correct the answer ?
Answer #1
REDD COMPANY
Horizontal Statements Model - Year 2
Event Balance Sheet Income Statement Statement of Cash Flows
Assets = Liabilities + Stockholders’ Equity Revenue Expenses = Net Income
Cash + Accounts Receivable + Merchandise Inventory = Accounts Payable + Common Stock + Retained Earnings
Bal. 8,000 + + 2,000 = + 7,500 + 2,500 = NA
1a. + + 5500 = 5500 + + = 0 NA
1b. -500 + + 500 = + + = -500 OA
2 + + -350 = -350 + + = 0 NA
3a + + -103 = -103 + + 103 = NA
3b -5,047 = -5,047 -5047 OA
4a. + 9,000 + = + + 9,000 9,000 = 9,000 0 NA
4b. + + -6,000 = + + -6,000 6,000 = -6,000 0 NA
5a. -800 + + = + + -800 -800 = -800 -800 OA
5b. + + 500 = + + 500 -500 = 500 0 NA
6 -600 + + = + + -600 600 = -600 -600 OA
7a + -180 + = + + -180 -180 = -180 NA
7b 8820 -8,820 8820 OA
8 + + -347 = + + -347 347 = -347 0 NA
Bal. $9,873 + $0 + $1,700 = $0 + $7,500 + $4,176 $8,020 $6,447 = $1,573 $1,873 NA